Ace Ltd is a listed parent company with interests in television stations, cinemas and newspapers. On 1 January 2014, Ace Ltd acquired 40% of the voting shares of Deuce Pty Ltd, a publisher of women magazines, for $1 620 000 cash. The acquisition gave Ace Ltd. significant influence over Deuce Ltd. The recorded net assets and contingent liabilities of Deuce Ltd as at the date of acquisition were represented by the following equity items:

                                  $000
Share Capital             1,000
Retained Earnings       600
General Reserve          200
Total                            1,800


Additional information:

(a)   At the date of acquisition, Deuce Ltd has created several magazine mastheads. The terms and conditions of the mastheads indicate they can be transferred to another party. The costs relating to the development of these mastheads had been written off by Deuce Ltd as expenses when incurred. Ace Ltd can reliably measure the fair value of the unrecognised mastheads at the date of acquisition at $300 000.

(b)   Ace Ltd has adopted an accounting policy for the Ace Ltd extended group whereby all intangible assets with a finite life are to be amortised on a straight-line basis over their useful lives. Ace Ltd expects the mastheads will provide future economic benefits for a period of 20 years.

(c)   During the year ended 31 December 2014 Deuce Ltd earned profit before tax of $900 000, incurred an income tax expense of $300 000 and paid a dividend of $100 000 on 30 September 2014.

(d)   On 1 July 2014 Deuce Ltd sold Ace Ltd a printing machine at an agreed value of $420 000. This equipment had a carrying amount of $120 000 to Deuce Ltd at the date of its transfer. The remaining useful life of the machine at the date of transfer is estimated to be 3 years.

(e)   Ace Ltd uses the cost method to account for its investment in Deuce Ltd in its separate financial statements as there is no quoted market price for Deuce Ltd. shares.

(f)   Ace Ltd has not recognised any impairment losses in relation to its investment in Deuce Ltd in its separate financial statements or its consolidated financial statements for the year ended 31 December 2014.

(g)   The company tax rate is 30%.

Required:

i)   Calculate the amount of goodwill on acquisition of Act Ltd’s interest in Deuce Ltd and related journal entry under cost method.

(ii)   Prepare the equity accounting consolidation adjusting entries required in Ace Ltd’s consolidated financial statements for the year ended 31 December 2014.

(iii)   Estimate the carrying value of Ace Ltd’s investment in Deuce Ltd the year ended 31
Answer #1

i) Calculation of goodwill on acquisition of Act Ltd's interest in Deuce Ltd:

Goodwill amount = Consideration amount - Value of net assets taken over

Consideration = $1,620,000

Recorded net assets of Deuce Ltd.    $1,800,000
Add: Fair value of mastheads $300,000
Total $2,100,000
Value of net assets taken over $840,000
(40% of above)

Thus, value of goodwill = $(1,620,000-840,000) = $780,000

Journal entry:

Investment in Deuce Ltd. A/c Dr. $840,000
Goodwill A/c Dr. $780,000
To Cash Alc. $1,620,000

ii) Equity accounting consolidation adjusting entries required in Ace Ltd’s consolidated financial statements for the year ended 31st Dec, 2014:

Value of printing machine on transfer date = $120,000

Annual Depreciation of Printing Machine = $120,000/3 = $40,000

Depreciation till sale to Ace Ltd's = $40,000*6/12 = $20,000

Therefore, Profit on sale for Deuce Ltd = $420,000 - ($1,20,000 - $20,000)

= $320,000

Profit share for Ace Ltd's = $320,000 * 40%

= $128,000

Equity accounting consolidation adjusting entries as on 31st Dec, 2014:

Investments A/c Dr. $240,000

To Profit & Loss A/c $240,000

Profit & Loss A/c Dr. $128,000
To Investment  Alc. $128,000

iii) Carrying value of Ace Ltd's investment in Deuce Ltd for the year ended 31 December 2014:

Value of mastheads = $300,000
Depreciation for the year = Total value/Estimated Useful Life

= 300,000/20 = $15,000

Adjusted Profit share in Deuce Ltd = (Profit - Income Tax Expense) * 40% - Profit Share on sale of printing machine

= ($900,000 - $300,000) * 40% - $128,000

= $240,000 - $128,000

= 112,000

Depreciation related to Mastheads for Ace Ltd = $15,000 * 40% = $6,000
Total carrying value of Ace Ltd's investments in Deuce Ltd = Purchase Value - Depreciation related to Mastheads for Ace Ltd + Adjusted Profit share in Deuce Ltd

= $16,20,000 - ($15,000 * 40%) + $112,000

= $16,20,000 - $6,000 + $112,000

= $17,26,000